McCullough Hospital uses a job-order costing system to assign costs to its patients. Its direct materials include a variety of items such as pharmaceutical drugs, heart valves, artificial hips, and pacemakers. Its direct labor costs (e.g., surgeons, anesthesiologists, radiologists, and nurses) associated with specific surgical procedures and tests are traced to individual patients. All other costs, such as depreciation of medical equipment, insurance, utilities, incidental medical supplies, and the labor costs associated with around-the-clock monitoring of patients are treated as overhead costs.

Historically, McCullough has used one predetermined overhead rate based on the number of patient-days (each night that a patient spends in the hospital counts as one patient-day) to allocate overhead costs to patients. Recently a member of the hospital’s accounting staff has suggested using two predetermined overhead rates (allocated based on the number of patient-days) to improve the accuracy of the costs allocated to patients. The first overhead rate would include all overhead costs within the Intensive Care Unit (ICU) and the second overhead rate would include all Other overhead costs. Information pertaining to the hospital’s estimated number of patient-days, its estimated overhead costs, and two of its patients—Patient A and Patient B—is provided below:

ICU Other Total
Estimated number of patient-days 3,400 27,200 30,600
Estimated fixed overhead cost $ 5,375,400 $ 21,678,400 $ 27,053,800
Estimated variable overhead cost per patient-day $ 291 $ 112
Patient A Patient B
Direct materials $ 5,900 $ 7,600
Direct labor $ 28,500 $ 38,800
Total number of patient-days (including ICU) 23 28
Number of patient-days spent in ICU 0 16

Required:

1. Assuming McCullough uses only one predetermined overhead rate, calculate:

a. The predetermined overhead rate.

b. The total cost, including direct materials, direct labor and applied overhead, assigned to Patient A and Patient B.

2. Assuming McCullough calculates two overhead rates as recommended by the staff accountant, calculate:

a. The ICU and Other overhead rates.

b. The total cost, including direct materials, direct labor and applied overhead, assigned to Patient A and Patient B.
Answer #1

Answers

  • Requirement 1

--Part [a]

ICU

Other

Total

A

Variable overhead per PDs

$291

$112

B

Estimated no of PDs

3400

27200

30600

C = A x B

Total Variable overhead

$989,400

$3,046,400

$4,035,800

D

Total Fixed overhead

$5,375,400

$21,678,400

$27,053,800

E = C+D

Total Overhead

$31,089,600

F

Total no of patient days (PDs)

30600

G = E/F

Predetermined overhead rate

$1,016

--Part [b]

Patient A

Patient B

Direct Material

$5,900

$7,600

Direct Labor

$28,500

$38,800

Manufacturing overhead assigned

$23,368

$28,448

Total Cost

$57,768

$74,848

--Working

Patient A

Patient B

Direct Material

5900

7600

Direct Labor

28500

38800

Manufacturing overhead assigned

=23*1016

=28*1016

  • Requirement 2

--Part [a]

ICU

Other

A

Variable overhead per PDs

$291

$112

B

Estimated no of PDs

3400

27200

C = A x B

Total Variable overhead

$989,400

$3,046,400

D

Total Fixed overhead

$5,375,400

$21,678,400

E = C+D

Total Overhead

$6,364,800

$24,724,800

F = E/B

Overhead rates = ANSWER

$1,872

$909

--Part [b]

Patient A

Patient B

Direct Material

$5,900

$7,600

Direct Labor

$28,500

$38,800

Manufacturing overhead assigned

$43,056

$40,860

Total Cost

$77,456

$87,260

--Working

Patient A

Patient B

Direct Material

5900

7600

Direct Labor

28500

38800

Manufacturing overhead assigned

=23*1872

=+(16*1872)+((28-16)*909)